Which is the solution to |3x - 2| = 6?
Answer by asuar010(338)

to solve absolute value equations we just set up the first equation as 3x-2=6 which leads 3x=8 and x=8/3 as the first answer and the second equation as 3x-2=-6 which leads to 3x=-4 which leads to x=-4/3.